### v2

# Air-Ground Collaborative Robots for Fire and Rescue Missions: Towards Mapping and Navigation Perspective

Abstract
Air-ground collaborative robots have shown great potential in the field of fire and rescue. Mapping and navigation, as the key foundation for air-ground collaborative robots to achieve efficient task execution, have attracted a great deal of attention. This growing interest in collaborative robot mapping and navigation is conducive to improving the intelligence of fire and rescue task execution, but there has been no comprehensive investigation of this field to highlight their strengths. In this paper, we present a systematic review of the ground-to-ground cooperative robots for fire and rescue from a new perspective of mapping and navigation. First, an air-ground collaborative robots framework for fire and rescue missions based on unmanned aerial vehicle (UAV) mapping and unmanned ground vehicle (UGV) navigation is introduced. Then, the research progress of mapping and navigation under this framework is systematically summarized, including UAV mapping, UAV/UGV co-localization, and UGV navigation, with their main achievements and limitations. Based on the needs of fire and rescue missions, the collaborative robots with different numbers of UAVs and UGVs are classified, and their practicality in fire and rescue tasks is elaborated, with a focus on the discussion of their merits and demerits. In addition, the application examples of air-ground collaborative robots in various firefighting and rescue scenarios are given. Finally, this paper emphasizes the current challenges and potential research opportunities, rounding up references for practitioners and researchers willing to engage in this vibrant area of air-ground collaborative robots.
Index Terms:
Air-ground collaborative robots, fire and rescue robots, mapping, localization, navigation, path planning.
I Introduction
In daily life, various emergencies such as fires may occur at any time, threatening people’s lives and property safety. Fire and rescue, as a vital public safety service, bears irreplaceable responsibilities [1, 2, 3]. When encountering a fire incident, firefighters need to first conduct a quick and accurate assessment of the scene to determine the intensity and the spread speed of the fire, as well as the areas that may be affected. They then rationalize their efforts to extinguish the fire and search for trapped persons according to the fire situation.
With the development of robotics and artificial intelligence technology, air-ground collaborative robots based on unmanned aerial vehicle (UAV) and unmanned ground vehicle (UGV) have been proven to be effective in fire and rescue missions [4, 5, 6, 7]. It can provide more information support and operational flexibility for fire and rescue, which is of great significance to improving the efficiency and intelligence of search and rescue missions [8]. For instance, in fire scene investigation, UAVs can provide high-altitude perspectives and global observations to help firefighters understand the spread of the fire, and the location of the fire source, etc., while UGVs can search on the nearby ground to provide close-range fire monitoring and data collection, especially for complex terrain or areas that require detailed searches. In addition, since the operating costs of UAVs and UGVs are relatively low, the air-ground collaborative robots can effectively save the use of search and rescue resources, as well as reduce the risks and possibility of injuries to rescuers [9, 10]. To this end, air-ground collaborative robots have the advantages of UAV and UGV, and have been widely used in fire and rescue missions through mutual cooperation.
For air-ground collaborative robots, mapping and navigation technology, as the key foundation for realizing efficient and autonomous collaborative working between UAV and UGV, has received increasing amounts of attention [11, 12, 13, 14, 15]. In general, with the help of the superior mobility of UAV to quickly perceive large-scale environmental information, on the one hand, a coarse-grained full-scene map for UGV navigation is established, and on the other hand, the mission scene is efficiently obtained [16]. Based on this, the UGV can achieve rapid path planning and safe navigation to perform fire and rescue missions [8, 17]. During this period, refined modeling can be achieved in combination with the local information perceived by the UGV to provide more comprehensive environmental information. Therefore, the collaborative working paradigm of air-ground robots is a promising solution to improve the efficiency, accuracy and safety of fire and rescue missions.
Numerous review and survey articles related to air-ground collaborative robots have been presented. Chen et al. [18] presented a taxonomy of UAV and UGV cooperative systems and a generalized optimization framework, with the goal of showing how to achieve a common mission goal and perform optimization for specific tasks. Chai et al. [19] discussed the integration of UAV and UGV, focusing on their synchronous planning and control mechanisms. Ding et al. [20] reviewed the research progress of air-ground collaborative system and identified four key functional roles. Besides, Duan et al. [21] focused on the core issues of heterogeneous collaboration between UAVs and UGVs, and discussed heterogeneous clustering, formation control, formation stability, network control, and practical applications. Liu et al. [12] introduced the basic elements of UAV/UGV collaboration and classified its task types, metrics, and scenarios. Tang et al. [22] reviewed the use of swarm intelligence algorithms in multiple UAVs collaborations. Liu et al. [23] investigated the research results of UAVs and UGVs collaboration from the aspects of network design, resource allocation, performance analysis and optimization. Guo et al. [24] studied the service collaboration of air-ground collaborative robots to ensure the service quality of service computing. Similarly, Sheng et al. [25] explored the development and application examples of network technology based on UAVs and UGVs collaboration in the context of high-speed railways. Zhen et al. [26] discussed the research progress of air-ground collaborative mobile edge computing. Fei et al. [27] focused on the sensing and communication networks of air-ground collaborative robots and introduced the recent progress of system architecture and protocols.
Different from the above existing survey literature, this paper reviews the mapping and navigation methods of air-ground collaborative robots for fire and rescue missions from a new perspective, which is guided by the requirements of UAVs/UGVs to perform fire and rescue operations in a collaborative manner. First, a system architecture of air-ground collaborative robots for fire and rescue is introduced. Then, within this framework, we discuss the different types of maps built by UAV for UGV navigation, focusing on their applicability for UGV navigation. On this basis, the path planning and navigation methods of UGV are outlined from the perspective of different map types. Furthermore, for fire and rescue tasks, the air-ground collaborative robots are divided into four categories based on the number of UAVs and UGVs, and their different characteristics are elaborated. In addition, the case analysis on the application of air-ground collaborative robots in fire and rescue scenarios is provided. Finally, the existing challenges and future research directions are discussed. The contributions of this paper mainly include:
  1. 1.
    system architecture of air-ground collaborative robots for fire and rescue, as well as the discussion of the characteristics and applicability of mapping and navigation methods under this framework;
  2. 2.
    classification and case analysis of air-ground collaborative robots in fire and rescue missions;
  3. 3.
    elucidation on the main research challenges and suggestions for potential research opportunities in the future.
The remainder of this work is organized as follows. Section II introduces the air-ground collaborative navigation framework based on the requirements of performing fire and rescue missions. In Section III, the different types of maps constructed by UAVs for UGV navigation are discussed, which is the basisn. Section IV summarizes the navigation methods of UGVs from the perspectives of positioning and path planning, while Section V classifies and discusses the collaborative navigation system according to the different numbers of UAVs/UGVs. In Section VI, the application of mapping and navigation in fire and rescue scenarios is presented, and specific cases are explained. Finally, the conclusion of the paper and the trends for future research of air-ground collaborative mapping and navigation in the field of fire and rescue are given in Section VII.
II System architecture of air-ground collaborative robots
Air-ground collaborative robots play an increasingly irreplaceable role in fire and rescue, which can effectively improve the efficiency and safety of fire and rescue operations. As shown in Fig. 1, once a fire occurs, air-ground collaborative robots can quickly enter the fire scenario to perceive the location, scale and environmental conditions of the fire, and perform fire and rescue missions. According to mission requirements, environment mapping and navigation are the key foundations for achieving this goal. Typically, UAVs and need to be able to quickly build environmental maps thanks to their large-scale environmental perception capabilities, assisting UGVs in planning and navigation in challenging environments [28].
Figure 1: Illustration of air-ground collaborative robots framework for fire and rescue missions based on UAV mapping and UGV navigation.
In general, UAVs are equipped with advanced sensors such as high-definition cameras, lidars, and thermal imagers [29], which can quickly obtain real-time situation and heat distribution of disaster areas. With these data, UAVs can build different forms of maps according to the scene characteristics and mission requirements, providing important reference information for UGVs to navigate safely and efficiently [30, 31]. Based on this, UGV plans an optimal path by considering information such as terrain and fire-fighting mission points, and achieves navigation with minimal risks to complete the fire and rescue task [32, 33]. For complex and large-scale rescue scenarios, air-ground collaborative robots can be expanded to multiple UAVs and UGVs to work together to improve the efficiency. For example, multiple UAVs can perceive, map, and monitor different areas respectively, and which is then is exchanged in real time for coordination and decision-making. Further, according to the task assignment, multiple UGVs use the information perceived by the UAVs to navigate to the assigned task points to collaboratively perform firefighting and rescue tasks.
In the above collaborative tasks, this paper mainly focuses on the map construction and navigation of the air-ground collaborative robots, and the schematic of their relationship is shown in Fig. 2. The UAVs are used for rapid environmental perception to build maps, and realize the cooperative localization of the UAVs and UGVs. On this basis, the UGVs perform path planning and navigation with the help of the map constructed by UAVs to complete fire and rescue missions.
Figure 2: Schematic of the relationship between air-ground collaborative robots for map construction and navigation.
III UAV Map Building for UGV Navigation
Map construction is a prerequisite for robots to complete fire and rescue tasks, which provides a basis for robotic motion planning and navigation. In air-ground collaborative operations, thanks to the superiority of UAV in perception and motion, map construction is usually completed by UAV, and then used for UGV navigation. To meet the needs of UGV navigation, the maps constructed by UAV are typically divided into 2-D grid map, 3-D map, topological map, and semantic map. According to the characteristics of different scenarios, suitable map types can be selected to promote UGV navigation and efficient collaborative operations. The representative works are also summarized in Table I.
TABLE I: Summary of representative works on UAV map building for UGV navigation
Map Type Reference Feature Experimental Evaluation
Robots (UAV) Sensor Scenario
2-D Grid Map Chatziparaschis et al. [34] Constructing a 2-D map, integrated with 3D OctoMap, to produce a navigable map for UGVs DJI Matrice 100 Hokuyo UTM-30LX-EW LiDAR, Stereo camera University campus
Katikaridis et al. [35] Inserting the constructed map into ROS as a Portable Graymap (PGM) image format for efficient UGV navigation DJI Spreading Wings S1000+ Sony RX100 III RGB camera Agricultural environments
Li et al. [36] Using the occupancy map built by UAV, the perception-limited UGVs conduct path planning for navigation Quadrotor Livox Mid-360 LiDAR Real-world simulation scenarios
3-D Map Delmerico et al. [37] Using UAV for active exploration and employing monocular 3-D reconstruction from low-altitude to map the terrain Quadrotor MAV Downward-looking camera, IMU Driveway and canyon (Outdoor simulation)
Arbanas et al. [38] Exploring unknown environments, constructing 3-D maps by mapping obstacles and walls AscTec NEO hexacopter Skybotix VI-Sensor Narrow corridors, Maze
Forster et al. [39] Performing monocular dense 3-D reconstruction, using aerial views to enhance ground maps NanoQuad MAV Down-looking camera, IMU Indoor and outdoor environment
Topological Map Wang et al. [40] Constructing 3-D ESDF from an aerial perspective, generating a lightweight 2-D topological map - Intel realsense D400 Computer room and Shed / Simulation environment
Wang et al. [41] UAV autonomous exploration while constructing environmental topological maps Developed UAV platform Intel RealSense RGB-D camera Indoor environment
Gomez et al. [42] Using the Free-area graph and the Topological map to represent information - Hokuyo laser sensor, Asus Xtion camera Simulation environment / Real-world office
Semantic Map Yue et al. [43] Constructing a semantic map through a hierarchical cooperative probabilistic semantic mapping framework UAV (with Intel NUC 6i7KYK @ 2.60 GHz CPU) ZED stereo camera Open Carpark, Indoor-Outdoor mixed environment
Bultmann et al. [44] Integrating multi-modal semantic fusion and label propagation techniques to build a semantic map in real-time DJI Matrice 210 v2 3D-LiDAR, RealSense D455 Urban environment, Disaster test site
Miller et al. [45] Using cell-wise semantic object categories as a common map representation for UAV and UGV teams The Falcon 4 quadrotor Global shutter RGB camera, ZED-F9P GPS Outdoor environment
III-A 2-D Grid Map
The image data obtained by the UAV can be converted into a 2-D grid map through image processing and object recognition technology [46, 47]. The grid map divides the area into grid cells, each of which represents a different type of object or obstacle information. Such map can be used for UGV path planning, obstacle avoidance and navigation decisions [48].
Li et al. [49] constructed a ground map by processing ground images acquired by UAV using image denoising and correction techniques. The map is then used for UGV path planning and navigation. Chatziparaschis et al. [34] designed a ground-air robot collaboration method, in which the UAV establishes an occupancy grid map by extracting the area related to the UGV, and then the UGV performs the path planning based on the created map for safe navigation. Similarly, Katikaridis et al. [35] investigated a UAV-supported UGV path planning system. The system uses environmental information obtained by the UAV as input, which was then converted into a grid-based map after preprocessing to support UGV path planning. In addition, Li et al. [36] proposed an air-ground collaborative architecture to solve the autonomous navigation problem of a group of blind UGVs. In this architecture, the UAV is used for environmental perception and grid map construction to guide a group of UGVs to navigate in an unknown environment (refer to Fig. 3). For robot collaborative operations, the 2-D grid map constructed by UAG has the characteristics of simplicity, easy to generate and update, and it is suitable for UGV navigation and path planning, with small data volume and computing requirements [50, 51]. However, for complex operation scenarios, the 2-D map representation capability is limited and cannot provide detailed perceptual information such as height and semantics, which makes it difficult for UGVs to adapt to complex terrain.
Figure 3: Experimental scenario of UAV-guided UGVs navigation, where the UAV provides environment perception and mapping for UGVs to avoid potential collisions [36].
III-B 3-D Map
3-D maps provide UGVs with a stereoscopic view of the ground surface, including information such as terrain, buildings, trees, and other height features [52, 53, 54]. Typically, 3-D maps are represented in the form of point cloud maps, octree maps, etc. These maps can be used for high-precision positioning, obstacle avoidance, and path planning for UGVs, and are particularly suitable for complex environments or rugged terrain.
For example, Delmerico et al. [37] proposed an air-ground collaborative active exploration method that uses a UAV to perceive and generate a 3-D map of the environment to provide the UGV with traversable paths through unknown terrain. This method can effectively shorten the response time for providing assistance in real disaster scenarios. Arbanas et al. [38] utilized a UAV to explore the environment, aiming to establish a 3-D map in the form of an OctoMap (see Fig. 4). This map is then projected onto the ground to guide the UGVs through the environment. Similarly, Wang et al. [55] developed a ground-air collaborative navigation system that combines an octomap-based 3-D environment mapping method for map creation in complex environments. Moreover, Fedorenko et al. [56] constructed a 3-D point cloud of the environment using a UAV and then investigated the feasibility of using this point cloud map directly for UGV planning and navigation. In the case that the initial relative attitude of UAV and UGV is unknown, Pei et al. [57] only used a laser scanner as a sensor to realize ground-air cooperative point cloud fusion and 3-D map construction. For air-ground collaboration, Forster et al. [39] used a monocular camera carried by a UAV to construct a dense 3D point cloud map to enhance the efficiency of UGV positioning and navigation. Hence, 3-D maps can not only provide rich 3-D information, but also provide a more realistic and accurate representation of the geographical environment to support air-ground collaborative robots to operate effectively in complex environments [58, 59]. However, the data volume of 3-D maps is large, and the processing and storage requirements are high. Also, the data acquisition and updating are more complicated than 2-D maps.
Figure 4: Scenario of the parcel delivery task with UGV based on OctoMap built by UAV [38].
III-C Topological Map
As an abstract map representation, topological maps are usually represented by nodes with edges, where nodes represent key locations in the environment, and edges represent direct or indirect connections between nodes [60]. For air-ground collaborative robots, topological maps can simplify road networks [61]. In this context, UAVs can generate road network maps based on topological relationships. Such maps describe the connection relationships and topological structures of roads for UGV path planning and navigation.
Wang et al. [40] proposed a method for constructing a two-dimensional topological map of the ground environment from an aerial perspective using UAVs. This method can be effectively applied to global path planning for UGVs. Wang et al. [41] used the depth camera carried by the UAV to build a topological map that is completed incrementally as the UAV flies, as shown in Fig. 5. This map explicitly shows the topological structure of the 3-D environment, and can also effectively provide information gain and cost-to-go for the candidate areas to be explored. To improve the efficiency of global planning, Blochliger et al. [62] put forward Topomap, a framework for simplifying navigation tasks, by converting sparse feature maps into 3-D topological maps. Gomez et al. [42] developed a hybrid mapping method that combines the efficiency of topological maps with the accuracy of dense representations for path planning. Besides, Zhang et al. [7] proposed a cross-domain collaborative unmanned framework called GACF. This framework utilizes a topological construction algorithm based on aerial perspective images from UAVs for road network extraction, providing a global traversable map for UGV path planning purposes. Such topological map constructed with UAVs provides a simplified road network structure, which can efficiently support UGVs’ path planning and navigation decisions on the basis of reducing data volume and computing requirements [63]. It is more suitable for areas with relatively simple road network structures such as cities and highways [64]. However, due to the lack of detailed geometric information and environmental characteristics, it has certain limitations in complex environments such as fire and rescue scenarios.
Figure 5: A topological map of the 3-D environment that is constructed incrementally as the UAV flies within the sensor range [41].
III-D Semantic Map
Semantic maps can provide richer environmental perception and scene understanding for air-ground collaborative robots [65, 66]. In this context, UAVs can use visual sensors to build semantic maps containing geographical elements such as roads, buildings, and traffic signs to support high-level decision-making and behavior planning of UGVs.
Vasić et al. [67] performed semantic segmentation on the images captured from the UAV’s viewpoint using deep neural networks, thereby providing effective path planning for successful navigation of the UGV. Maturana et al. [68] proposed a semantic mapping system for automatic off-road driving of UGVs, which provides a richer environmental representation than pure geometric maps. This can provide insights for the navigation of UGVs in firefighting and rescue scenarios. Furthermore, Yue et al. [43] established a hierarchical collaborative probabilistic semantic mapping framework by combining information from UAVs and UGVs to obtain semantic point clouds, which are then utilized for generating semantic maps. Bultmann et al. [44] designed a multi-sensor model for real-time semantic reasoning and fusion of UAV systems (see Fig. 6). It is capable of fusing and generating a global semantic map, holding tremendous potential in fast autonomous or remote-controlled semantic scene analysis. This system can be applied to disaster inspections. Miller et al. [45] exploited an integrated air-ground collaboration system where a semantic map created in real-time by the UAV enables UGVs to perform localization, planning, and navigation. The system is demonstrated to be effective in simulating disaster scenarios. However, there is still room for improvement in the local planning phase. Semantic maps emphasize the semantics and attribute information of geographic features in maps. They can provide richer scene understanding and environmental perception to support advanced decision-making and behavior planning of UGVs, and are suitable for complex urban environments [69, 70]. However, semantic maps have limitations such as large data volume, high processing and storage requirements, and high creation and maintenance costs.
Figure 6: Semantic perception and mapping of disaster scenes [44].
IV UAV Map-Based UGV Navigation
In the navigation of air-ground collaborative robots, the core links include environment mapping, localization and path planning. For fire and rescue tasks, after the UAV creates the corresponding map according to the task, the UGV uses its own sensors to identify the surrounding environmental features based on these maps to achieve positioning. Subsequently, the UGV plans a safe and efficient route to ensure that the fire and rescue mission can be successfully completed. Map construction has been explained in Section III, and localization and path planning will be discussed below.
IV-A UAV and UGV Co-Localization
Co-Localization is an indispensable component of air-ground collaborative robot navigation, and its accuracy directly affects the success rate and efficiency of task execution. Typically, localization methods can be divided into GPS/GNSS localization, inertial navigation system (INS) localization, visual localization, lidar localization, and multi-sensor fusion localization, as shown in Table II.
TABLE II: Summary of Representative Works on UAV and UGV Co-Localization
Localization Methods Reference Merit Demerit
GPS/GNSS localization Li et al. [71] Positioning through information sharing and consistency checks Not delved deeply into solving the integrity algorithm issues for inter-vehicle and multi-sensor integration in multi-robotic systems
Correa-Caicedo et al. [72] Having some competitiveness in GPS data correction; With relatively low dependence on parameters and sensors Limited generalization capability, need retraining in new geographic regions to fit the data
INS localization Yilmaz et al. [73] Exploring for GPS-denied environments Lack of validation of robustness in real scenarios
Zhou et al. [74] Provide attitude information and offer continuous measurement data (INS) IMU measurements accumulate errors, lowering positioning accuracy
Visual localization Zhang et al. [75] Taking into account the unstructured and feature-deficient nature of the environment Insufficient features may lead to inaccurate positioning or failure
Song et al. [76] Combining semantic information for localization; Considering the relocalization efficiency in dynamic indoor environments Coarse positioning accuracy depends on detection precision
Lidar localization Lima et al. [77] Fusing aerial and ground perception data; Exploring the positioning issue in outdoor unstructured environments without GPS Lack of evaluation of the impact of environmental appearance changes on the robustness of positioning
Caballero and Merino [78] Based on raw point cloud data; Evaluating the robustness under odometry errors Low spatial efficiency; Large memory required
Multi-sensor fusion localization Kayhani et al. [79] Based on camera and IMU; Considering the cost-effectiveness and lightweight design of the localization method Need to manually place or replace the label in the scene
Patoliya et al. [80] Integrating GPS and LiDAR data; Considering error accumulation in an unbounded environment Limited in mapping large open space; Insufficient consideration for use on non-planar surfaces
Yousuf et al. [81] Integrating INS, GPS, and odometer data; Having a certain impact on positioning accuracy in indoor and outdoor environments Lack of consideration for sensor noise and error in real environments
IV-A1 GPS/GNSS localization
The GPS or GNSS can provide basic coordinate information for UAVs and UGVs for rough localization [82, 83]. Based on this, collaborative robots use the localization information in the same coordinate system to perform coordinate transformation to allow them to use a common map for navigation. Li et al. [71] designed a GNSS-based collaborative localization method that uses differential pseudo-range and carrier phase observations to obtain the accurate relative position between robots, as shown in Fig. 7. In addition, the GPS-based localization solution is very sensitive to the environment in which the robot is located, and it may experience system failures due to environmental conditions. To solve this problem, Correa-Caicedo et al. [72] proposed an intelligent system based on fuzzy logic that obtains information from sensors and corrects the absolute position of the robot based on longitude and latitude. GPS can improve the accuracy and reliability of localization and has a wide coverage range. GPS-based localization strategies are suitable for outdoor open environments, but may not work properly indoors or in other places where GPS signals are poor, and are easily affected by obstructions [84].
Figure 7: Top view of the trajectory drawn on Google Earth, as well as the available satellite numbers of the UGV [71].
IV-A2 INS localization
INS measures the acceleration of the carrier in the inertial reference frame and integrates this data over time. And it is then converted into the navigation coordinate system, so that the information such as the velocity, yaw angle and position of the carrier in the navigation coordinate system can be obtained to achieve robot localization [85, 86]. Yilmaz et al. [73] designed an INS-based robot localization method, in which INS provides reference coordinates and uses the angle and linear position of the Denavit-Hartenberg method to calculate the robot’s position based on the reference INS module. Traditional INS-based localization has poor accuracy and severe drift due to error accumulation. To solve this problem, Zhou et al. [74] proposed a tightly integrated localization system based on Wi-Fi round-trip time, encoder, and inertial measurement unit (IMU). Through the error state Kalman filter, the inertial information from the IMU is fused with the measurement results from the encoder to suppress the accumulated error of INS. Attentively, INS does not rely on external signals. Moreover, INS can work in a variety of environments such as air, ground, and underwater, and is not restricted by weather and lighting conditions [87, 88]. However, INS also has some limitations. For example, INS is very sensitive to vibration and temperature changes, which may affect its accuracy, and in complex scenes, the alignment time is long.
IV-A3 Visual localization
For air-ground collaborative robots, the vision-based localization solution is usually that the UAV uses visual sensors to collect environmental data to build a scene map, and then the UGV obtains information about the surrounding environment to match the scene map and database to achieve the localization of the UGV [89, 90]. Zhang et al. [75] suggested a visual localization method that adopts fiducial markers and factor graphs to achieve reliable localization performance. In this method, fiducial markers are introduced into the front-end module, and then the detection algorithm AprilTag2 is used to identify the markers and extract feature points to locate the robot. The visual localization method can solve the problem of localization and navigation of air-ground collaborative systems without GPS [91]. It can model scene maps based only on visual sensors, and realize robot localization through feature matching and map registration [76]. It can effectively reduce system costs. However, the performance of visual localization will deteriorate when the light is insufficient or the environment changes greatly, and it usually requires strong processor support.
IV-A4 Lidar localization
Lidar localization is to align the geometric feature data collected by the lidar sensors with the map of the surrounding environment for air-ground collaborative robot localization [77, 92]. Caballero and Merino [78] introduced a direct lidar localization scheme. This method uses a point cloud to map registration based on nonlinear optimization of the distance between the points and the 3-D map for robot localization (see Fig. 8). Such solution can achieve robot localization and navigation in areas without GPS, but the memory required to build the 3-D map is very large. To this end, Zhang et al. [93] designed a dual-lidar high-precision natural navigation system based on the ROS platform. The system adopts the Lidar-SLAM method based on graph optimization to construct a 2-D environmental map, and then uses the PF algorithm in MRPT for robot localization. Lidar-based technology is very effective in detecting obstacles and accurately measuring distances [94]. It is not affected by lighting conditions and is suitable for use in a variety of scenarios [95]. However, the cost of lidar is relatively high, and in some cases it may be affected by bad weather such as dust and rain.
Figure 8: UAV firefighting scenario and constructed 3D map with lidar data for localization [78].
IV-A5 Multi-sensor fusion localization
The multi-sensor fusion localization method can effectively address the defects of a single sensor, enabling the robot to achieve more accurate localization and navigation in different environments and conditions [96, 97]. Donati et al. [98] designed a method based on GPS, IMU and ultrasound sensors. The method combines the data provided by multiple sensors with the information provided by 3-D low-complexity maps to support the localization of air-ground collaborative robots. Kayhani et al. [79] proposed a tag-based visual inertial localization method, which takes advantage of the vision to identify the tag position and uses the manifold EKF to fuse inertial data and tag measurements for UAV localization. In addition, Patoliya et al. [80]exploited a robot localization system that integrates GPS and LiDAR data. Similarly, Yousuf et al. [81]introduced a method for robot localization that fuses INS, GPS and odometer sensors. The key to multi-sensor fusion localization technology lies in how to effectively process and analyze data from different sensors, and how to maintain the accuracy and stability of localization in a dynamically changing environment [99]. With the continuous advancement of sensor technology, fusion localization is becoming more and more widely used in the field of robotics [100, 101], which helps to improve the localization performance and reliability of robots in various application scenarios.
IV-B UGV Navigation
UGV efficient navigation is a key step for air-ground collaborative robots to complete firefighting and rescue missions. To achieve this, UGV needs to be able to select the appropriate navigation method based on the map constructed by the UAV. In such context, according to the type of map established by the UAV, the UGV navigation method is classified into navigation based on 2-D grid map, 3-D map, topological map, and semantic map. A summary of the representative works is also presented in Table III.
TABLE III: Summary of Representative Works on UGV Navigation Based on UAV Maps
Applied Map Types Reference Feature Experimental Evaluation
Sensor Path Planning Method Scenario
2-D grid map-based Zhang et al. [102] Based on the occupancy grid map, combined with virtual areas, performing global and local path planning for human-aware indoor navigation SICK TIM561 laser rangefinder D* algorithm / DWA Indoor environment
Zhang et al. [103] Constructing a 2-D grid map and using Bidirectional RRT with assisting metric for path planning to achieve indoor navigation Laser Radar Bi-AM-RRT* Indoor environment
3-D map-based Delmerico et al. [37] Using UAV monocular 3-D reconstruction to create topographic maps, UGVs calculate paths for navigation Rangefinder, Camera, IMU D* algorithm (Modified version) Driveway and Canyon (Outdoor simulation)
Arbanas et al. [38] Utilizing the map constructed by UAV for path planning, UGV accomplishes navigation through the cooperation of UAV VI-Sensor Lattice-based algorithm Narrow corridors, Maze
Topological map-based Wang et al. [40] Using a 2-D topological map constructed by UAVs, UGVs perform global path planning for navigation Lidar, Camera A* search Computer room and Shed / Simulation environment
Blochliger et al. [62] Constructing topological maps based on sparse visual features, generating paths with a lightweight topological planner to achieve navigation VI sensor, Laser A* search Semistructured industrial site
Ravankar et al. [104] Building the topological map on the grid map, using the A* algorithm for topological navigation Hokuyo lidar A-star algorithm Simulated environment
Zuo et al. [105] Constructing reduced approximated generalized Voronoi graph, employing a three-phase rapid path search method for 3-D space navigation Velodyne VLP-16 laser scanner Three-stage grid-graph-combined approach / Dijkstra Underground parking, Warehouse and Treelawn, Factory
Semantic map-based Kostavelis et al. [106] Integrating semantic mapping methods with robotic navigation strategies to construct semantic maps and achieve hierarchical navigation Camera Dijkstra algorithm COLD dataset (Ljubljana section)
Deng et al. [107] Integrating semantic segmentation and SLAM frontend to generate dense semantic maps, using semantic information to enhance path planning algorithms Realsense D435 camera Algorithm incorporating semantic information RRL competition environment
Vasić et al. [67] Using the semantic map constructed from the UAV’s perspective for UGV path planning in dynamic environments to achieve navigation Lidar, Camera A variant of the Dijkstra algorithm Car parks
IV-B1 2-D grid map-based navigation
2-D grid maps can be directly used for UGV navigation, allowing the UGV to plan the optimal path from the current position to the target position [102]. Representatively, Dijkstra [108] proposed a Dijkstra algorithm for shortest path, which solves the single-source shortest path problem of weighted directed or undirected graphs by using breadth-first search. In order to improve the operation speed of Dijkstra, Hart et al. [109] developed an A* algorithm, which uses a heuristic function to predict the target position, thereby giving priority to exploring areas that appear to be closer to the target. This allows the robot to find the target in fewer steps. In response to the problems of long calculation time and many search nodes in A*-based global path planning, a variety of improved algorithms based on search and sampling are proposed for UGV planning and navigation (refer to Fig. 9) [110, 103, 111, 112]. Although the UGV navigation method based on 2-D grid can achieve efficient and safe navigation, its navigation performance is subject to the accuracy of the map, and it has limitations such as poor adaptability to 3-D obstacles and highly dynamic scenes.
Figure 9: An example of 2-D map-based path planning and navigation [110].
IV-B2 3-D map-based navigation
3-D maps are usually not used directly for UGV navigation. They are usually converted into 2-D or 2.5-D maps for navigation. Among them, 2.5-D maps are most suitable for navigation [113]. This is because they can not only use 2-D map navigation algorithms to find feasible paths for UGVs, but also represent the height of the terrain. For example, Delmerico et al. [37] combined the altitude and terrain data of the 3-D map and used the Dijkstra algorithm to plan a navigable path for the UGV that takes into account both the difficulty of the terrain and the shortest completion time, thereby improving the navigation efficiency in complex environments, as shown in Fig. 10. This is particularly important for tasks such as exploration, firefighting and rescue. Analogously, Basso et al. [114] designed an air-ground collaborative 3-D map sharing strategy, in which the UAV can contribute the constructed 3-D map to the UGVs for navigation. Arbanas et al. [38] taken advantage of 3-D point cloud maps for terrain assessment and directly employed the rapidly explored random tree (RRT) algorithm on the point cloud map for effective path planning. The advantage of navigation based on 3-D maps is that it can provide deeper environmental information for UGVs to choose from [114], while the disadvantage is that the computing resources and costs required for UGVs to process and update these complex data are relatively high.
Figure 10: 3-D map including elevation information, and the feasible path planned by UGV [37].
IV-B3 Topological map-based navigation
UGV can directly use the topological map constructed by UAV to plan the reference line for navigation, which can effectively reduce the path planning time [60, 115]. Blochliger et al. [62] proposed Topomap, a framework for creating multifunctional topological maps and reliable navigation, which uses a graph search algorithm to traverse topological points to obtain the shortest path. Wang et al. [40] used a graph search algorithm to search for a global path based on the 2-D topological map drawn by the UAV, and obtained a safe and feasible global path to achieve autonomous navigation of the UGV. Oleynikova et al. [116] constructed a sparse graph containing 3-D topological information for robot planning and navigation in a 3-D environment, as shown in Fig. 11. Ravankar et al. [104] proposed a hybrid topological mapping and navigation method for mobile robots, using the Djikstra and A* algorithms to perform navigation between nodes. Zuo et al. [105] introduced the Voronoi diagram into the topological mapping model. On this basis, a three-stage fast path search and navigation method was designed. Topological maps allow UGVs to perform efficient path planning and navigation. However, due to the large amount of debris in fire and rescue scenarios, it is difficult for topological maps to accurately draw the obstacles and feasible paths, making it challenging for UGVs to safely navigate based on topological maps.
Figure 11: Sparse 3-D topological maps constructed by UAVs that can be used for UGVs navigation in 3-D environments [116].
IV-B4 Semantic map-based navigation
Semantic maps can provide UGVs with richer environmental information (e.g., scene and object semantics or traversable and intraversable areas), so that UGVs can better perceive and understand the surrounding environment [117]. Kostavelis et al. [106] constructed a multi-layer semantic map considering the concepts of time and space, and provided robots with a means of layered navigation solutions. Deng et al. [107] designed a semantic SLAM framework for rescue robot navigation. With the help of semantic information, rescue robots can identify different types of terrain in complex environments to select safe navigation routes with better traversability (refer to Fig. 12). Vasić et al. [67] segmented the traversable areas of images acquired by UAVs based on deep neural networks, and proposed a UGV path planning method with semantic maps in dynamic environments. Semantic maps can support UGVs in semantic planning and navigation. However, the creation and updating of semantic maps require a lot of computing resources [118]. In addition, for complex and dynamic fire and rescue environments, it is extremely challenging to accurately identify and segment objects, areas and other information, which will affect the modeling quality of semantic maps and thus the navigation performance of UGVs.
Figure 12: Example of semantic maps including eight types of terrain for UGV navigation [107].
V Classification of Air-Ground Collaborative Robots in Fire and Rescue Missions
For fire and rescue missions, the number of UAVs and UGVs used in the air-ground collaborative robot system can vary depending on task requirements, environmental conditions, and available resources. In such context, the composition of air-ground coordination systems can generally be classified into four categories: a single UAV with a single UGV, a single UAV with multi-UGVs, multi-UAVs with a single UGV, and multi-UAVs with multi-UGVs, as illustrated in Table IV.
TABLE IV: Summary of Representative Works on Air-Ground Collaborative Systems with different numbers of UAVs and UGVs
Classification Reference Merit Demerit
A single UAV with a single UGV Giakoumidis et al. [119] Considering the complementary advantages of UAV and UGV; Enhancing the UGV environmental perception capabilities and navigation flexibility Not yet expanded to include a greater number and types of robots; Not yet been tested in a full-scale outdoor environment
Fankhauser et al. [120] Utilizing the complementary advantages of UAV and ground robot; Having certain advantages in positioning and navigation in complex and unknown terrain Challenges in achieving reliable wireless communication; Yet to be validated in a broader and more complex environment
Mueggler et al. [121] Verified in simulated disaster scenarios; Having certain advantages in terms of autonomy Two main assumptions do not hold in reality; Not yet covered the collaboration of multiple aerial and ground robots
A single UAV with multi-UGVs Miller et al. [45] UGVs navigate using real-time semantic maps from UAV; Validity tested in simulated disaster scenarios Efficiency and scalability challenges exist for deploying and managing large-scale UAV-UGV teams
Yulong et al. [122] Demonstrating the speed and reliability of UAV collaborating with multiple UGVs in information gathering and transmission Low real-time computing performance
Ding et al. [123] Considering information transmission and collaborative operations across a large expanse No multi-UAVs collaboration considered
Multi-UAVs with a single UGV Stampa et al. [124] More comprehensively mapping the area; Showing some advantages in rescue efficiency Further study required for technology integration, environmental adaptability, and reliability
Ren et al. [125] Making certain progress on the path planning issue for multi-UAVs and UGV systems Complex path planning in collaborative systems
Peng et al. [126] Combining vehicle transport capacity with multi-UAVs task efficienc Need to consider UAVs path planning and task allocation issues
Multi-UAVs with multi-UGVs Qin et al. [127] Combining the agility of UAVs with the computational resources of UGVs; Optimizing perception and mapping in GPS-denied environments Multi-UAVs and multi-UGVs systems face communication, coordination, and task allocation challenges
Ghamry et al. [128] Making some progress in forest monitoring and fire detection with UAVs-UGVs teams Relatively simple UAVs formation reconstruction strategy
Phan et al. [129] Integrating the advantages of different types of vehicles; Considering the potential application in firefighting missions Insufficient exploration in algorithm performance
V-A A single UAV with a single UGV
For small-scale fire monitoring and rescue missions, the collaboration of a single UAV with a single UGV is widely used. In such a collaborative system, the UAV provides aerial surveillance and builds a map of the environment, while the UGV navigates based on the constructed map to perform specific tasks. Compared with the collaboration of multiple UAVs and UGVs, the tasks of the UAV and UGV in this collaboration are clearly divided, and the implementation of the localization and navigation is relatively simple.
Giakoumidis et al. [119] designed a heterogeneous architecture consisting of a UGV and a UAV, where the UAV serves as a separable remote vision system for the UGV. In the system, the environmental image information perceived by the UAV is spliced to construct a 2-D map, and then the FM2 algorithm is used to implement UGV path planning and collaborative navigation. This collaborative framework can be used for fire and rescue missions in unknown environments. Similarly, Fankhauser et al. [120] developed a collaborative navigation framework for flying and walking robots. The method uses the UAV’s onboard monocular camera to create a visual feature map for positioning and dense environment representation. Based on the established environmental model, the UGV positioning is realized and navigation is planned by interpreting the established map in terms of traversability. In order to enable air-ground collaborative robots to perform autonomous collaborative tasks in disaster scenarios, Mueggler et al. [121] taken advantage of the UAV to capture ground images for 2-D grid map construction, and employed the Dijkstra algorithm for UGV path planning. However, each action of the UGV is iteratively directed by the UAV, resulting in low autonomy. In addition, Christie et al. [130] proposed a UAV-UGV cooperative solution for task execution using scene understanding, as shown in Fig. 13. In this solution, a UAV is used to identify points of interest and model digital elevation of the unknown scene, and build 3-D semantic map based on semantic segmentation. Then based on these outputs, the UGV then utilizes the carried liDAR for collision-free path planning and autonomous task execution. Although this type of a single UAV with a single UGV collaborative operation has been successfully applied, it still has limitations such as low task execution capability and poor adaptability to complex environments for large-scale and complex fire and rescue scenarios [131].
Figure 13: A pipeline for UAV-UGV cooperative solution based on semantic segmentation [130].
V-B A single UAV with multi-UGVs
The collaboration of a UAV with multi-UGVs can improve the efficiency of mission execution and provide faster response capabilities. In this configuration, the UAV can provide comprehensive environmental information to UGVs, allowing UGVs to collaboratively perform search and rescue missions [132]. In addition, UGVs can also share map data, sensor information, etc. to improve collaborative work efficiency.
Miller et al. [45] proposed a semantic-based air-ground collaborative robot system framework that can realize collaborative navigation of a UAV and multiple UGVs. The system deploys a UAV to create a semantic map of the ground in real time, which is then used for UGVs’ positioning and semantic navigation tasks. To improve the navigation efficiency of multi-UGVs missions, Zhang et al. [133] suggested a strategy to organize multi-UGVs formations. In this system, a frontier-based exploration strategy is used to search for cooperative targets on the map, and the A* algorithm is then employed to plan the collision-free path through all targets. Similarly, Zang et al. [134] designed a multi-UGVs system solution based on coordinated behavior planning and trajectory planning. However, the coordination problem between UAV and multiple UGVs is not considered. In addition, in the fire and rescue missions, UGVs may be scattered over a large area, which may affect their collaboration due to the limited perception capabilities. In this case, UAV not only needs to perform map construction to assist UGV navigation, but also needs to act as a messenger to support indirect communication of UGV. Yulong et al. [122] addressed the problem of path planning for a messenger UAV in a system involving aerial and ground collaboration with multiple UGVs. This provides a solution for achieving one-to-many air-ground collaborative navigation, facilitating real-time information sharing and coordinated rescue operations in fire and rescue scenarios. To solve the path planning problem of collaborative heterogeneous robots consisting of a UAV and multiple UGVs, Ding et al. [123] developed an effective memetic algorithm to find the shortest route that allows the messenger UAV to access all moving UGVs. This treatment allows the UGVs to perform fire and rescue missions in a collective manner over a large area (refer to Fig. 14), and improve mission efficiency to a certain extent. However, for large-scale and complex fire and rescue scenarios, this kind of a single UAV with multi-UGVs collaboration requires UAV to have long-term perception and environmental modeling capabilities, which in turn makes the single UAV face challenges such as long flight time.
Figure 14: A scenario for a single UAV with multi-UGVs performing fire and rescue tasks in a large area [123].
V-C Multi-UAVs with a single UGV
The configuration of multi-UAVs with a single UGV is suitable for fire and rescue missions that require large-scale monitoring and data collection. Multiple UAVs work together to provide a more comprehensive perspective and perception capabilities, which can complete map construction faster to support UGV navigation and task execution. In addition, multiple UAVs also increase the fault tolerance and robustness of the collaborative system to a certain extent.
Stampa et al. [124] designed an air-ground collaborative system consisting of multiple UAVs and a UGV as a mobile base station, as depicted in Fig. 15. In this system, the data collected by the UAVs are fused together, and then a 3-D map of the environment is constructed based on SLAM technology to assist the navigation of robots and human rescue missions. When the accurate environmental map is completed, the system switches to monitoring mode, with some UAVs providing a real-time aerial overview of the scene and quickly detecting environmental changes to effectively address the complexity and dynamics of the fire and rescue environment. Scaramuzza et al. [135] described a collaborative system consisting of multiple UAVs that can achieve autonomous navigation in large, unknown and GPS-free environments. These works mainly involve map construction of multi-UAV systems, but pay less attention to UGV navigation based on maps. For the path planning problem in multi-UAVs-UGV collaborative mode, Ren et al. [125] developed a framework for path planning in an air-ground collaborative system consisting of a pair of UAVs and a UGV. This architecture can be used for tasks such as collaborative search and rescue of heterogeneous robotic systems. Regarding the collaborative operation of UGV-UAV, Peng et al. [126] gave a routing and scheduling algorithm for multiple UAVs and a UGV, which allows multiple UAVs carried by the UGV to perform rescue tasks such as delivering multiple packages in different locations at the same time. The above work supports the collaborative execution of firefighting and rescue missions by UAVs and a UGV in large-scale scenarios. Although multiple UAVs can provide efficient environmental perception and map construction, the efficiency and timeliness of a single UGV in executing such emergency tasks cannot be guaranteed.
Figure 15: Illustration of multi-UAVs with a single UGV performing fire and rescue tasks [124].
V-D Multi-UAVs with multi-UGVs
Multiple UAVs with multiple UGVs are suitable for complex and large-scale fire and rescue missions, and can provide all-round perception and efficient mission execution. It not only has the perception advantages of multiple UAVs, such as rapid deployment over the fire area to achieve efficient search and map building, and the ability to quickly adjust strategies to continue to guide UGV operations when some UAVs fail; but also has the load-bearing and mission execution advantages of multiple UGVs, such as simultaneous search and rescue work at different locations, as well as the delivery of rescue supplies.
Qin et al. [127] designed a collaborative exploration and mapping method for UAVs and UGVs. The method uses UGVs to generate a rough environmental map based on 2.5-D SLAM algorithm, and then the UAVs perform 3-D fine mapping to support task execution. Ghamry et al. [128] developed a multi-UAVs and multi-UGVs collaborative system to monitor forests and detect fires. When the UAV detects a fire, it immediately transmits the relevant fire information to the leader UGV. After receiving the fire information, the leader UGV determines the best route quickly based on path planning algorithm, and guides other nearby UGVs to the fire scene through navigation. At the same time, the UAV will track the UGV to detect and extinguish the spread of the fire. Nazarova et al. [136] proposed applying the air-ground collaborative system to search and rescue missions to improve search and rescue efficiency, as illustrated in Fig. 16. First, the sequence of rescue actions and rescue processes is analyzed based on statistical data. Then, based on the search theory of probability theory and the use of intelligent optimization algorithms, the air-ground collaborative system is ensured to complete the task more efficiently. In order to achieve efficient collaboration between multiple UAVs and multiple UGVs, Phan et al. [129] suggested a hierarchical architecture including mobile mission controller, multiple UAVs and multiple UGVs, in which the UAVs are responsible for collecting environmental data and assisting in fire monitoring, while the UGVs are used to extinguish large open fires. Tanner [137] combined the decentralized flocking method with the navigation function to develop an exchange cooperative control scheme to coordinate UAVs and UGVs to locate mobile targets in a given area. Through the close collaboration of multiple UAVs and multiple UGVs, the efficiency of mission execution is significantly improved, ensuring rapid response and efficient management of rescue operations [19].
Figure 16: Illustration of a multi-UAVs with multi-UGVs collaborative rescue system, in which UAVs are used for environmental perception and search for survivors, and UGVs are adopted to provide emergency rescue [136].
VI Application examples of air-ground collaborative robots
Air-ground collaborative robots have demonstrated their unique advantages in firefighting and rescue missions, among which collaborative mapping and navigation technology plays an important role. Such collaborative operation mode can achieve rapid response and efficient handling of complex environments such as forest wildfires, urban high-rise building fires, mountain rescue, chemical plant fires, etc., to improve the efficiency and safety of firefighting and rescue operations.
At present, air-ground collaborative robots have been implemented in practical applications and test experiments in various scenarios. Frering et al. [5] provided a collaborative robots system based on UAV and UGV to assist firefighting and rescue missions, as shown in Fig. 17. In the system, the UAV is equipped with RGB and thermal infrared cameras to detect hot spots and conduct aerial reconnaissance. At the same time, images and temperature data of the fire scene are obtained and a 3-D map of the fire scene is constructed. The UGV can extinguish the fire source with water based on the map information. The system is field tested in the Seetaler Alps in Austria and demonstrates outstanding performance and reliability.
Figure 17: Collaborative robots perform firefighting missions. Top: execution process. Bottom: visualization on the client [5].
In order to deal with forest fires and rescue tasks, Wei and Fang [138] proposed a collaborative framework based on UAVs and UGVs, as presented in Fig. 18. The UAVs use a coverage path planning algorithm to comprehensively monitor the fire scene and guide the evacuation planning of UGVs, while The UGVs adopt a reactive algorithm to plan a collision-free path to collaboratively complete the evacuation task. This solution can effectively improve the efficiency and safety of forest fire and rescue tasks.
Figure 18: Collaborative framework for forest fires and rescue tasks. (a):Collaborative framework. (b): UAVs cruising and mapping. (C): UGV path planning [138].
For wildfire prevention in outdoor scenarios, Pasini et al. [139] proposed a wildfire hotspot surveillance system based on UAV and UGV cooperation. The system uses the UAV to monitor and mapping the area after the fire from the air, and then uses the A* algorithm to plan a path for the UGV. The UGV conducts centimeter-level inspections along the planned path and employs a short-term correction algorithm to flexibly avoid unexpected obstacles. The system provides a low-cost and efficient solution for wildfire hotspot surveillance.
Figure 19: Path planning and navigation results of UAV and UGV cooperation for wildfire hotspot surveillance [139].
For firefighting and rescue scenarios, Stampa et al. [124] designed a air-ground collaborative system (refer to Fig. 20(a)). In the system, the UAV uses RGB-D to build a 3-D map of the environment and detect fires. The UGV performs path planning and navigation based on the map built by the UAV and drags a hose to extinguish the fire, which improves the efficiency of firefighting and reduces the risk to personnel. Similarly, Martinez-Rozas et al. [140] introduced a UAV/UGV collaborative team for autonomous firefighting in urban environments, which consists of three UAVs and a UGV. The UAVs take advantage of 3-D LIDAR for mapping and use infrared thermal imagers to sense the location of the fire. The UGV utilizes a combination of global path planning and local path tracking for navigation. This system is verified in the high-rise fire rescue scenarios, as illustrated in Fig. 20(b).
Figure 20: Air-ground collaborative robots for firefighting and rescue scenarios. (a): Conceptual illustration of the air-ground collaborative system [124]. (b): Application scenario of high-rise building fire and rescue mission [140].
In summary, air-ground collaborative robots have achieved remarkable results in fire and rescue. They use UAVs to perform environmental perception and mapping, and guide UGVs navigation to efficiently complete fire and rescue tasks. This not only greatly improves the efficiency of firefighting and rescue, but also significantly reduces the risks directly faced by operators.
VII Conclusion and Outlook
The mapping and navigation technology of air-ground collaborative robots is the core foundation for realizing the efficient and autonomous collaboration of UAVs and UGVs to complete fire and rescue tasks. To this end, guided by the needs of performing fire and rescue tasks, this paper reviews the mapping and navigation methods of air-ground collaborative robots for fire and rescue tasks from a new perspective. First, we introduce an air-ground collaborative navigation framework in which UAVs perform mapping and UGVs implement navigation on this basis, and emphasize its importance in improving rescue efficiency and personnel safety. Then, this paper discusses the mapping methods of different UAV map types that can be used for UGV navigation, including 2-D grid map, 3-D map, topological map, and semantic map, as well as their characteristics and applicability. Besides, the process of UGV navigation, including co-Localization and path planning, is explained, focusing on different Localization methods, and reviewing commonly used path planning methods from the perspective of different map types. Furthermore, according to the number of UAVs/UGVs, the collaborative robot systems are classified, and the advantages and disadvantages of a single UAV with a single UGV, a single UAV with multi-UGVs, multi-UAVs with a single UGV, and multi-UAVs with multi-UGVs collaborative schemes are discussed. Finally, the application examples of air-ground collaborative robots in fire rescue scenarios are given to demonstrate the applicability of the collaborative robots. This paper systematically introduces the air-ground collaborative navigation method for fire rescue, providing a valuable reference for further improving rescue efficiency and personnel safety in fire rescue missions. To sum up, this paper can provide a valuable reference for how to use air-ground collaborative robots in fire rescue missions to improve rescue efficiency and personnel safety.
Although the air-ground collaborative robots have shown great success in the field of fire rescue, there are still challenges in using UAVs to build efficient maps that are conducive to UGV navigation and task execution, as well as high-precision navigation and efficient task execution of UGVs for complex fire and rescue scenarios. For fire and rescue missions, there may be following potential research opportunities for air-ground collaborative robot mapping and navigation.
1) Fire and rescue missions-oriented UAV mapping. The existing maps, whether 2-D grid map, 3-D map, topological map, and semantic map, or other combined forms of maps, mostly focus on the map itself, while ignoring the relevance to the fire and rescue missions. Establishing an environment model that highlights task-related environmental features, semantics and other information has been proven to significantly improve the efficiency of robotic task execution [141, 142]. Therefore, combining the operating environment and task characteristics, it is a promising direction to use task point detection, semantic segmentation, heat map analysis, 3-D reconstruction, 2-D mapping, multi-UAV collaborative mapping and other technologies to build a map that is conducive to UGV navigation and task execution, and can provide information such as fire distribution, hotspot location and building structure to assist in the formulation of rescue strategies.
2) Multi-modal data fusion for UAV mapping. Fire and rescue environments are often complex and changeable, and are filled with smoke, flames, obstructions, etc., which pose challenges to UAV mapping. To address such challenges, multi-sensor SLAM systems have demonstrated high accuracy and robustness in complex environments [143, 144], and can effectively overcome the impact of smoke and lighting changes on visual perception [145]. To this end, the information of multiple sensors such as thermal imaging, visible light, and infrared can be integrated to perform multi-modal data fusion and UAV mapping. Such data fusion can effectively enhance the perception and mapping capabilities of UAVs, and also provide a more comprehensive description of the environment for UGVs. Therefore, the UAV mapping system with multi-modal data fusion will be an important research direction for air-ground collaborative robots in the fire and rescue environments.
3) UGV navigation based on UAV maps. Based on different types of maps built by UAVs, it is particularly important to study the general navigation of UGVs, which is the basis for realizing the execution of air-ground collaborative robot tasks. With the continuous integration of deep learning technology, the localization and perception capabilities of UGV in unknown or unstructured environments are enhanced [146, 147]. On this basis, by learning environmental features in real time, UGVs can achieve fast path planning and efficient obstacle avoidance based on different types of maps, further enhancing the adaptability of UGV navigation. Note that the improvement of edge computing capabilities can effectively improve the efficiency and real-time performance of UGV navigation based on deep learning, which will be a direction worth exploring for air-ground collaborative robots.
4) Embodied artificial intelligence (AI)-driven UGV navigation. The combination of embodied AI and robotics technology has brought profound impact on robot navigation [148]. And some work has verified that the integration with embedded AI, especially large language models (LLMs) and visual language models (VLMs), significantly enhances the robot’s scene understanding, path planning and autonomous navigation capabilities in complex environments [149, 150, 151]. In fire and rescue missions, the combination of collaborative navigation systems and embodied AI can, on the one hand, improve the performance of UAV/UGV real-time fire scene understanding, task area identification and feature matching for UGV navigation, and on the other hand, effectively enhance the navigation performance and mission execution capabilities of UGV. This will play an integral role in the research of air-ground cooperative robot navigation.